# LeanAgent: Lifelong Learning for Formal Theorem Proving

### v6

@@ Abstract
Large Language Models (LLMs) have been successful in mathematical reasoning tasks such as formal theorem proving when integrated with interactive proof assistants like Lean. Existing approaches involve training or fine-tuning an LLM on a specific dataset to perform well on particular domains, such as undergraduate-level mathematics. These methods struggle with generalizability to advanced mathematics. A fundamental limitation is that these approaches operate on static domains, failing to capture how mathematicians often work across multiple domains and projects simultaneously or cyclically. We present LeanAgent, a novel lifelong learning framework for theorem proving that continuously generalizes to and improves on ever-expanding mathematical knowledge without forgetting previously learned knowledge. LeanAgent introduces several key innovations, including a curriculum learning strategy that optimizes the learning trajectory in terms of mathematical difficulty, a dynamic database for efficient management of evolving mathematical knowledge, and progressive training to balance stability and plasticity. LeanAgent successfully proves 162 theorems previously unproved by humans across 23 diverse Lean repositories, many from advanced mathematics. It performs significantly better than the static LLM baseline, proving challenging theorems in domains like abstract algebra and algebraic topology while showcasing a clear progression of learning from basic concepts to advanced topics. In addition, we analyze LeanAgent’s superior performance on key lifelong learning metrics. LeanAgent achieves exceptional scores in stability and backward transfer, where learning new tasks improves performance on previously learned tasks. This emphasizes LeanAgent’s continuous generalizability and improvement, explaining its superior theorem-proving performance.
@@ 1 Introduction
Figure 1: LeanAgent overview. LeanAgent searches for Lean repositories and uses LeanDojo to extract theorems and proofs. It uses curriculum learning, computing theorem complexity as $e^{S}$ ($S$ = proof steps), and calculating the 33rd and 67th complexity percentiles across all theorems to sort repositories by easy theorem count. LeanAgent adds the curriculum to its dynamic database. As a retrieval-based framework, LeanAgent generates a dataset (premise corpus and a collection of theorems and proofs) for each repository in the curriculum and progressively trains its retriever. Progressive training happens over one epoch to prevent forgetting old knowledge. Then, LeanAgent uses the updated retriever in a search-based method to prove results previously unproved by humans, known as sorry theorems. It adds new proofs to the database.
@@
LeanAgent. We present LeanAgent, a novel lifelong learning framework for theorem proving. As shown in Figure 1, LeanAgent’s workflow consists of (1) deriving complexity measures of theorems to compute a curriculum for learning, (2) progressive training to learn while balancing stability and plasticity, and (3) searching for proofs of sorry theorems (theorems unproven by humans) by leveraging a best-first tree search. LeanAgent works with any LLM; we implement it with retrieval for improved generalizability [81]. LeanAgent incorporates several key innovations. It uses a custom dynamic database to manage its evolving mathematical knowledge. LeanAgent uses a novel curriculum learning strategy utilizing the structure of Lean proofs to learn on increasingly complex mathematical repositories.
@@
Extensive experiments across 23 diverse Lean repositories demonstrate LeanAgent’s advancements in lifelong learning for theorem proving. LeanAgent successfully proves 162 theorems unproven by humans, known as sorry theorems, many from advanced mathematics. This count includes mathematical proofs and exploits found within Lean’s type system. For example, it proves difficult sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology related to Coxeter systems and the Hairy Ball Theorem [18, 30]. We find that LeanAgent demonstrates progressive learning in theorem proving, initially proving basic sorry theorems and significantly advancing to more complex ones. It significantly outperforms the static ReProver baseline in terms of proving new sorry theorems while maintaining performance on known ones. We have issued pull requests to the respective repositories with the newly proven sorry theorems. Some of them turned out to be exploits due to the loopholes in the repositories that are being fixed, and in future, should no longer be an issue in the future after the fixes. We plan to run LeanAgent periodically on all the repositories as they are being updated with human-written proofs, enabling LeanAgent to continue learning and prove more theorems over time.
@@ 3 Methodology
A useful lifelong learning strategy for theorem proving requires (a) the best repository order strategy and (b) the best learning strategy. We solve (a) with curriculum learning to utilize the structure of Lean proofs and (b) with progressive training to balance stability and plasticity. LeanAgent consists of four main components: curriculum learning, dynamic database management, progressive training of the retriever, and sorry theorem proving. Further methodology details are in Appendix A.1 and a discussion of why curriculum learning works in theorem proving is available in Appendix A.6.
@@ 3.3 Progressive Training of the Retriever
Although LeanAgent works with any LLM, we provide a specific implementation here. We start with ReProver’s retriever, a fine-tuned version of the ByT5 encoder [79], leveraging its general pre-trained knowledge from mathlib4. We train LeanAgent on the new dataset for an additional epoch. This limited exposure helps prevent overfitting to the new data while allowing LeanAgent to learn essential new information. Before validation, we precompute embeddings for all premises in the corpus to ensure these embeddings are consistent with LeanAgent’s current state. To understand how LeanAgent balances stability and plasticity, we save the model iteration with the highest validation recall for the top ten retrieved premises (R@10). This is a raw plasticity value: it can be used to compute other metrics that describe LeanAgent’s ability to adapt to and handle new types of mathematics in the latest repository (details in Sec. 4). Then, we compute the average test R@10 over all previous datasets the model has progressively trained on, a raw stability value.
@@ 4.1 Experimental Setup
[…]
Figure 2: Case studies of LeanAgent’s new proofs. LeanAgent shows an ability to work with these repositories, often able to retrieve the necessary premises (highlighted). For example, LeanAgent proves a sorry theorem from PFR, condRho_of_translate, by simply expanding definitions, showing its proving ability on the PFR repository. It proved a PFR theorem with just rfl (by reflexivity), demonstrating its grasp of how the terms in the theorem statement interrelate. In addition, LeanAgent could prove re_float from SciLean during lifelong learning, while ReProver could not. Moreover, its MiniF2F proofs demonstrate its ability to generate relatively longer and more complex proofs for complex mathematics. LeanAgent’s proof of wedderburn and invmap.of_eq represents its theorem proving capabilities with abstract algebra premises.
Table 2: Theorem Proving Performance Score (TPPS) across repositories. Repositories with 0 sorry theorems or LeanAgent TPPS are not shown. The best TPPS for each repository is in bold.
Repository LeanAgent ReProver
PFR 11 1
Hairy Ball Theorem 11 1
Coxeter 11 1
PFR (alternate commit) 12 2
Mathematics in Lean Source 85 15
Formal Book 13 3
MiniF2F 226 86
SciLean 55 25
Carleson 2 2
Lean4 PDL 2 2
In addition, a use case of LeanAgent is formalizing in a new repository after learning a curriculum; we progressively train on MiniF2F to demonstrate this. Note that we choose the Lean4 version of the MiniF2F repository [80] and disregard its separation into validation and test splits (reasoning in Appendix A.5). Moreover, a mathematician could use LeanAgent for (1) an initial curriculum $A$, and later (2) a sub-curriculum $B$. LeanAgent can then help the mathematician formalize the repositories in curriculum $A+B$ in parallel. To demonstrate this scenario, we continue LeanAgent on a sub-curriculum $B$ of 8 repositories.
@@
LeanAgent demonstrates continuous generalizability and improvement in theorem-proving capabilities across multiple repositories. By the end of lifelong learning, LeanAgent has an 11$\times$ improvement factor over ReProver on PFR, 5.67$\times$ on Mathematics in Lean Source, 2.63$\times$ on MiniF2F, and 2.2$\times$ on SciLean. Additionally, LeanAgent had an 11$\times$ improvement factor over ReProver on the Hairy Ball Theorem, 11$\times$ on Coxeter, and 4.33$\times$ on Formal Book. Its proofs are a superset of the sorry theorems proved by ReProver in most cases. LeanAgent progresses from basic concepts (arithmetic, simple algebra) to advanced topics (abstract algebra, topology).
PFR. LeanAgent can prove a sorry theorem from this cutting-edge repository, while ReProver cannot. It also generalizes to a different commit (not included in progressive training), proving a theorem with just the rfl tactic that ReProver cannot. Interestingly, LeanAgent has a strong enough ability of logical manipulation within the PFR repository to prove 5 sorry theorems with a $0=1$ placeholder theorem statement (not included in TPPS) (details in Appendix A.5).
SciLean. During lifelong learning, LeanAgent proves theorems related to fundamental algebraic structures, linear and affine maps, and measure theory basics. By the end of lifelong learning, it masters concepts in advanced function spaces, sophisticated bijections, and abstract algebraic structures.
Mathematics in Lean Source. During lifelong learning, LeanAgent proves theorems about basic algebraic structures and fundamental arithmetic properties. By the end of lifelong learning, it proves complex theorems involving quantifier manipulation, set theory, and relations.
MiniF2F. ReProver demonstrates proficiency in basic arithmetic, elementary algebra, and simple calculus. However, by the end of lifelong learning, LeanAgent grasps advanced number theory, sophisticated algebra, complex calculus and analysis, abstract algebra, and complex induction.
Sub-curriculum. In the Formal Book repository, LeanAgent progresses from proving basic real analysis and number theory theorems to mastering advanced abstract algebra, exemplified by its proof of Wedderburn’s Little Theorem. For the Coxeter repository, LeanAgent proves a complex lemma about Coxeter systems, showcasing its proficiency in group theory. In the Hairy Ball Theorem repository, LeanAgent proves a key step of the theorem, demonstrating performance in algebraic topology. Only LeanAgent can prove these impressive theorems, demonstrating that it has much more advanced theorem-proving capabilities than ReProver.
Limitations of TPPS: There is no standardized measure of proof difficulty and proving performance, especially in a life-long learning setting. TPPS was an attempt towards this. The core idea is to have a higher reward for newly proven theorems, which aligns with the objective of lifelong learning. In future works, we also plan to investigate some more sophisticated measures that reward not only newly proved theorems but also difficult ones. However, TPPS has several important limitations. It is vulnerable to artificial inflation through exploitation of technicalities, does not account for varying theorem difficulties, and lacks standardization across mathematical domains. The metric focuses on quantity and faces challenges in ensuring comparisons across diverse repositories. These shortcomings highlight the need for more sophisticated evaluation frameworks in theorem proving, especially for lifelong learning contexts. Future work should focus on developing standardized, robust metrics that address these issues and provide more meaningful assessments of theorem-proving capabilities.
@@ 4.3 Lifelong Learning Analysis
WF5 0.18 7.60 7.17 0.73 15.83 2.23 13.34 5.82
FM 0.85 6.53 4.04 2.11 10.50 4.06 11.44 3.80
CFR 0.88 0.87 0.88 0.85 0.76 0.94 0.75 0.90
EBWT 1.21 0.51 1.04 0.76 -0.20 0.73 -1.34 -0.39
WP5 2.47 0.89 1.47 3.42 0.00 0.09 0.00 0.11
IP 1.02 0.36 0.26 1.06 -1.50 -0.64 -1.71 -0.89
CS 0.94 0.16 0.47 0.61 0.16 0.97 0.04 0.78
@@
Consistency with sorry Theorem Proving Performance. This lifelong learning analysis is consistent with LeanAgent’s sorry theorem proving performance. LeanAgent’s superior stability metrics (WF5, FM, and CFR) explain its ability to maintain performance across diverse mathematical domains, as evidenced by its success in proving theorems from various repositories like SciLean, Mathematics in Lean Source, and PFR. Its high EBWT score aligns with its progression from basic concepts to advanced topics in theorem proving. While LeanAgent shows slightly lower plasticity (WP5 and IP) compared to some setups, this trade-off results in better overall performance, as reflected in its ability to prove a superset of sorry theorems compared to other methods. The composite score, consisting of continuous generalizability, continuous improvement, and plasticity, corroborates LeanAgent’s overall superiority in lifelong learning for theorem proving.
@@ 5 Conclusion
We have presented LeanAgent, a lifelong learning framework for theorem proving that achieves continuous generalizability and improvement across diverse mathematical domains. Key components include a curriculum learning strategy, progressive training approach, and custom dynamic database infrastructure. LeanAgent proves 162 sorry theorems across 23 Lean repositories, including from challenging mathematics, highlighting its potential to assist in formalizing complex proofs across multiple domains and identifying system exploits. For example, LeanAgent successfully proves advanced sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology. It outperforms the ReProver baseline by up to 11$\times$, progressively learning from basic to highly complex mathematical concepts. Moreover, LeanAgent shows significant performance in forgetting measures and backward transfer, achieving a near-perfect composite score of 94%. This explains its continuous generalizability and continuous improvement.
Future work could explore integration with Lean Copilot, providing real-time assistance with a mathematician’s repositories. In addition, a limitation of LeanAgent is its inability to prove certain theorems due to a lack of data on specific topics, such as odeSolve.arg_x0.semiAdjoint_rule in SciLean about ODEs. To solve this problem, future work could explore using reinforcement learning for synthetic data generation during curriculum construction.

### v7

@@ Abstract
Large Language Models (LLMs) have been successful in mathematical reasoning tasks such as formal theorem proving when integrated with interactive proof assistants like Lean. Existing approaches involve training or fine-tuning an LLM on a specific dataset to perform well on particular domains, such as undergraduate-level mathematics. These methods struggle with generalizability to advanced mathematics. A fundamental limitation is that these approaches operate on static domains, failing to capture how mathematicians often work across multiple domains and projects simultaneously or cyclically. We present LeanAgent, a novel lifelong learning framework for formal theorem proving that continuously generalizes to and improves on ever-expanding mathematical knowledge without forgetting previously learned knowledge. LeanAgent introduces several key innovations, including a curriculum learning strategy that optimizes the learning trajectory in terms of mathematical difficulty, a dynamic database for efficient management of evolving mathematical knowledge, and progressive training to balance stability and plasticity. LeanAgent successfully proves 155 theorems previously unproved formally by humans across 23 diverse Lean repositories, many from advanced mathematics. It performs significantly better than the static LLM baseline, proving challenging theorems in domains like abstract algebra and algebraic topology while showcasing a clear progression of learning from basic concepts to advanced topics. In addition, we analyze LeanAgent’s superior performance on key lifelong learning metrics. LeanAgent achieves exceptional scores in stability and backward transfer, where learning new tasks improves performance on previously learned tasks. This emphasizes LeanAgent’s continuous generalizability and improvement, explaining its superior theorem-proving performance.
@@ 1 Introduction
Figure 1: LeanAgent overview. LeanAgent searches for Lean repositories and uses LeanDojo to extract theorems and proofs. It uses curriculum learning, computing theorem complexity as $e^{S}$ ($S$ = proof steps), and calculating the 33rd and 67th complexity percentiles across all theorems to sort repositories by easy theorem count. LeanAgent adds the curriculum to its dynamic database. As a retrieval-based framework, LeanAgent generates a dataset (premise corpus and a collection of theorems and proofs) for each repository in the curriculum and progressively trains its retriever. Progressive training happens over one epoch to prevent forgetting old knowledge. Then, LeanAgent uses the updated retriever in a search-based method to prove results previously unproved formally by humans, known as sorry theorems. It adds new proofs to the database.
@@
LeanAgent. We present LeanAgent, a novel lifelong learning framework for theorem proving. As shown in Figure 1, LeanAgent’s workflow consists of (1) deriving complexity measures of theorems to compute a curriculum for learning, (2) progressive training to learn while balancing stability and plasticity, and (3) searching for proofs of sorry theorems (theorems unproven formally by humans) by leveraging a best-first tree search. LeanAgent works with any LLM; we implement it with retrieval for improved generalizability [80]. LeanAgent incorporates several key innovations. It uses a custom dynamic database to manage its evolving mathematical knowledge. LeanAgent uses a novel curriculum learning strategy utilizing the structure of Lean proofs to learn on increasingly complex mathematical repositories.
@@
Extensive experiments across 23 diverse Lean repositories demonstrate LeanAgent’s advancements in lifelong learning for theorem proving. LeanAgent successfully proves 155 theorems unproven formally by humans, known as sorry theorems, many from advanced mathematics. For example, it proves challenging sorry theorems in abstract algebra and algebraic topology related to Coxeter systems and the Hairy Ball Theorem [17, 29]. LeanAgent also proves 7 theorems using exploits found within Lean’s type system. We find that LeanAgent demonstrates progressive learning in theorem proving, initially proving basic sorry theorems and significantly advancing to more complex ones. It significantly outperforms the static ReProver baseline in terms of proving new sorry theorems. We have issued pull requests to the respective repositories with the newly proven sorry theorems. Some of them turned out to be exploits due to the loopholes in the repositories that are being fixed, and in the future, should no longer be an issue after the fixes. We plan to run LeanAgent periodically on all the repositories as they are being updated with human-written proofs, enabling LeanAgent to continue learning and prove more theorems over time.
@@ 3 Methodology
A useful lifelong learning strategy for theorem proving requires (a) a repository order strategy and (b) a learning strategy. We solve (a) with curriculum learning to utilize the structure of Lean proofs and (b) with progressive training to balance stability and plasticity. LeanAgent consists of four main components: curriculum learning, dynamic database management, progressive training of the retriever, and sorry theorem proving. Further methodology details are in Appendix A.1 and a discussion of why curriculum learning works in theorem proving is available in Appendix A.6.
@@ 3.3 Progressive Training of the Retriever
Although LeanAgent works with any LLM, we provide a specific implementation here. We start with ReProver’s retriever, a fine-tuned version of the ByT5 encoder [78], leveraging its general pre-trained knowledge from mathlib4. We train LeanAgent on the new dataset for an additional epoch. This limited exposure helps prevent overfitting to the new data while allowing LeanAgent to learn essential new information. LeanAgent’s retriever, and therefore the embeddings it generates, are continuously updated during progressive training. Thus, at the end of the current progressive training run, we precompute embeddings for all premises in the corpus generated by LeanAgent’s current state to ensure that we properly evaluate LeanAgent’s validation performance. To understand how LeanAgent balances stability and plasticity, we save the model iteration with the highest validation recall for the top ten retrieved premises (R@10). This is a raw plasticity value: it can be used to compute other metrics that describe LeanAgent’s ability to adapt to and handle new types of mathematics in the latest repository (details in Sec. 4). Then, we compute the average test R@10 over all previous datasets the model has progressively trained on, a raw stability value.
@@ 4.1 Experimental Setup
We compare the number of sorry theorems LeanAgent can prove, both during and after lifelong learning, to the ReProver baseline. We use ReProver as the baseline because we use its retriever as LeanAgent’s initial retriever in our experiments.
Figure 2: Case studies of LeanAgent’s new proofs. LeanAgent shows an ability to work with these repositories, often able to retrieve the necessary premises (highlighted). For example, LeanAgent proves a sorry theorem from PFR, condRho_of_translate, by simply expanding definitions, showing its proving ability on the PFR repository. In addition, LeanAgent could prove re_float from SciLean during lifelong learning, while ReProver could not. Moreover, its MiniF2F proofs demonstrate its ability to generate relatively longer and more complex proofs for complex mathematics. LeanAgent’s proof of invmap.of_eq and wedderburn represents its theorem proving capabilities with abstract algebra premises.
Table 2: Number of proven sorry theorems across repositories. LA denotes LeanAgent. Note that “during” means the number of theorems proved during lifelong learning, “additional after” means the number of additional theorems proved after lifelong learning, and “total” is the number of theorems proved both during and after lifelong learning. “MIL” stands for Mathematics in Lean Source, “PFR (alternate)” refers to the alternate commit of PFR, and “Hairy Ball” refers to the Hairy Ball Theorem. “sorrys” denotes the total number of sorry theorems in each repository. Repositories with no sorry theorems or no proven ones are not shown. The best proving performance for each repository is in bold. As noted previously, we progressively train on MiniF2F after the initial curriculum to demonstrate the use case of formalizing in a new repository after learning a curriculum. As such, we don’t evaluate LeanAgent after lifelong learning on MiniF2F.
Repository sorrys LA (total) LA (during) LA (additional after) ReProver
MiniF2F 406 99 99 - 85
SciLean 294 27 22 5 24
MIL 29 21 14 7 14
Formal Book 29 3 2 1 2
PFR (alternate) 24 2 2 0 1
PFR 37 1 1 0 0
Hairy Ball 14 1 0 1 0
Coxeter 15 1 1 0 0
Carleson 24 1 1 0 1
Lean4 PDL 30 1 1 0 1
In addition, a use case of LeanAgent is proving in a new repository after learning a curriculum; we progressively train on MiniF2F to demonstrate this. Note that we choose the Lean4 version of the MiniF2F repository [79] and disregard its separation into validation and test splits (reasoning in Appendix A.5). LeanAgent’s success rate on the Lean4 version of the MiniF2F test set is also in Appendix A.5. Moreover, a mathematician could use LeanAgent for (1) an initial curriculum $A$, and later (2) a sub-curriculum $B$. LeanAgent can then help the mathematician prove in the repositories in curriculum $A+B$. To demonstrate this scenario, we continue LeanAgent on a sub-curriculum $B$ of 8 repositories.
@@
LeanAgent demonstrates continuous generalizability and improvement in theorem-proving capabilities across multiple repositories. LeanAgent’s proofs are a superset of the sorry theorems proved by ReProver in most cases. LeanAgent progresses from basic concepts (arithmetic, simple algebra) to advanced topics (abstract algebra, topology).
PFR. LeanAgent can prove a sorry theorem from this repository, while ReProver cannot. It also generalizes to a different commit (not included in progressive training), uncovering 7 system exploits. LeanAgent proves two theorems with just the rfl tactic, one of which ReProver cannot, and proves 5 sorry theorems with a $0=1$ placeholder theorem statement.
SciLean. During lifelong learning, LeanAgent proves theorems related to fundamental algebraic structures, linear and affine maps, and measure theory basics. By the end of lifelong learning, it proves concepts in advanced function spaces, sophisticated bijections, and abstract algebraic structures.
Mathematics in Lean Source. During lifelong learning, LeanAgent proves theorems about basic algebraic structures and fundamental arithmetic properties. By the end of lifelong learning, it proves more complex theorems involving quantifier manipulation, set theory, and relations.
MiniF2F. ReProver demonstrates proficiency in basic arithmetic, elementary algebra, and simple calculus. However, by the end of lifelong learning, LeanAgent handles theorems with advanced number theory, sophisticated algebra, complex calculus and analysis, abstract algebra, and complex induction.
Sub-curriculum. In the Formal Book repository, LeanAgent progresses from proving basic real analysis and number theory theorems to more advanced abstract algebra, exemplified by its proof of Wedderburn’s Little Theorem. For the Coxeter repository, LeanAgent proves a complex lemma about Coxeter systems, showcasing its increased understanding of group theory. In the Hairy Ball Theorem repository, LeanAgent proves a key step of the theorem, demonstrating improved performance in algebraic topology. Only LeanAgent can prove these impressive theorems, demonstrating that it has much more advanced theorem-proving capabilities than ReProver.
@@ 4.3 Lifelong Learning Analysis
WF5 ($\downarrow$) 0.18 7.60 7.17 0.73 15.83 2.23 13.34 5.82
FM ($\downarrow$) 0.85 6.53 4.04 2.11 10.50 4.06 11.44 3.80
CFR ($\uparrow$) 0.88 0.87 0.88 0.85 0.76 0.94 0.75 0.90
EBWT ($\uparrow$) 1.21 0.51 1.04 0.76 -0.20 0.73 -1.34 -0.39
WP5 ($\uparrow$) 2.47 0.89 1.47 3.42 0.00 0.09 0.00 0.11
IP ($\uparrow$) 1.02 0.36 0.26 1.06 -1.50 -0.64 -1.71 -0.89
CS ($\uparrow$) 0.94 0.16 0.47 0.61 0.16 0.97 0.04 0.78
@@
Consistency with sorry Theorem Proving Performance. This lifelong learning analysis is consistent with LeanAgent’s sorry theorem proving performance. LeanAgent’s superior stability metrics (WF5, FM, and CFR) explain its ability to maintain performance across diverse mathematical domains, as evidenced by its success in proving theorems from various repositories like SciLean, Mathematics in Lean Source, and PFR. Its high EBWT score aligns with its progression from basic concepts to advanced topics in theorem proving. While LeanAgent shows slightly lower plasticity (WP5 and IP) compared to some setups, this trade-off results in better overall performance, as reflected in its ability to prove a superset of sorry theorems compared to ReProver in most cases. The composite score, consisting of continuous generalizability, continuous improvement, and plasticity, corroborates LeanAgent’s overall superiority in lifelong learning for theorem proving.
@@ 5 Conclusion
We have presented LeanAgent, a lifelong learning framework for theorem proving that achieves continuous generalizability and improvement across diverse mathematical domains. Key components include a curriculum learning strategy, progressive training approach, and custom dynamic database infrastructure. LeanAgent proves 155 sorry theorems and uncovered 7 exploits across 23 Lean repositories, including from challenging mathematics, highlighting its potential to assist in formalizing complex proofs across multiple domains and identifying system exploits. For example, LeanAgent successfully proves challenging theorems in abstract algebra and algebraic topology. It outperforms the ReProver baseline in proving new sorry theorems, progressively learning from basic to complex mathematical concepts. Moreover, LeanAgent shows significant performance in forgetting measures and backward transfer, achieving a near-perfect composite score of 94%. This explains its continuous generalizability and continuous improvement.
Future work could explore integration with Lean Copilot, providing real-time assistance with a mathematician’s repositories. In addition, a limitation of LeanAgent is its inability to prove certain theorems due to a lack of data on specific topics, such as odeSolve.arg_x0.semiAdjoint_rule in SciLean about ODEs. To solve this problem, future work could explore using reinforcement learning for synthetic data generation during curriculum construction. Moreover, future work could explore using LeanAgent with additional math LLMs and search strategies.